OpenStudy (anonymous):

Help with variables on both sides problem? x-3/4=x/2

OpenStudy (anonymous):

Well here its not really that hard. Your problem looks like this? \[\frac{ x-3 }{ 4 }=\frac{ x }{ 2 }\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok well here you multiply criss-cross and set them equal to each other. Like this: 2(x-3)=4x Then solve!